# A Sober Look at Progress in Language Model Reasoning: Pitfalls and Paths to Reproducibility

### v1

@@ 4.4.1 Are Incorrect Responses Longer?
Yes. We find the trend is consistent across both RL- and SFT-trained models (additional figures provided in Appendix figures 17 and 18 ). We consistently observe that the effect is more pronounced in RL-trained models (displayed on the left) than in SFT-trained models (displayed on the right). As detailed in the Appendix, both the Qwen 2.5 Math base exhibit a slight shift in length, though this shift is notably more evident in R1-distill and subsequent RL-trained models.
@@
Although responses nearing the 32,000-token limit are almost always incorrect (due to limited context-length), this trend persists even for complete responses which are shorter– Longer responses are associated with a higher likelihood of being incorrect.
4.4.2 Is There Diversity Collapse in Reasoning Training?
AIME’24 | AIME’25 | AMC’23 | ||||||||
Model | Baseline | |||||||||
Open-RS3 | R1-Distill | -0.6 | +1.5 | -1.0 | +0.4 | +1.4 | -0.9 | -0.2 | +0.4 | |
DeepScaleR | R1-Distill | +0.6 | +0.2 | +3.6 | +2.2 | -1.8 | -1.7 | |||
S1.1-7B | Qwen-Instruct | +10.9 | +13.5 | +10.5 | +10.4 | +9.6 | +9.7 | |||
II-Thought | R1-Distill | -3.5 | -3.6 | +0.5 | +1.2 | +0.7 | +0.2 |
Dang et al. has reported a counterintuitive phenomenon in reasoning models: improvements in Pass@1 achieved through supervised fine-tuning or RL can reduce Pass@k performance due to diminished output diversity—a phenomenon termed diversity collapse. Theoretical analyses attribute this collapse to the model concentrating too much probability mass on a single reasoning path, while current decoding strategies fail to recover the lost diversity.
To examine these claims, we compare the Pass@k performance (for ) of RL-trained models against their corresponding base models (e.g., DeepSeek-R1-Distill-Qwen-1.5B) across all datasets. Table 4 shows the delta in Pass@k relative to each method’s base model.
Findings. We do not observe a consistent diversity collapse. Gains in Pass@1 generally come with improvements in Pass@k, though the magnitude of these gains varies. When Pass@k performance does drop, it does so alongside (rather than independently of) occasional declines in Pass@1, providing no support for the diversity collapse hypothesis.
@@ 5 Conclusion
Our study shows that much of the perceived progress in LLM-based reasoning, particularly in mathematical benchmarks, rests on unstable and often non-reproducible foundations. We find that minor differences in sampling parameters, prompt formatting, hardware, and software configurations can lead to major shifts in reported performance—casting doubt on many recent empirical claims. Reinforcement learning methods, while promising in theory, offer at best modest gains in practice and are prone to overfitting, especially on small benchmarks like AIME’24. In contrast, supervised finetuning continues to deliver consistent, generalizable improvements across a wide range of benchmarks and model sizes.
To address these challenges, we advocate for standardized, transparent evaluation protocols. Our open-sourced framework, complete with Dockerized environments, seed-averaged metrics, and robust answer matching, provides reproducible foundations for future research. We hope this work shifts the focus from leaderboard chasing to methodological rigor—ensuring that future claims of progress in reasoning are both meaningful and measurable.
@@ Acknowledgments
Appendix A Appendix
A.1 Bootstrapping Results on Additional Datasets
To complement our analysis in 3, we present bootstrapped variance results on two additional datasets: AMC’23 and MATH500. As shown in Figures 12 and 13, high variance in Pass@1 persists even when averaging over multiple seeds (), mirroring the trends observed on AIME’24. These results reinforce our conclusion that small benchmark sizes yield unstable estimates and that robust performance reporting requires multiple seed runs.
A.2 Hardware & Software Variations
In Figure 14, we show that the model performance variation due to hardware configuration is not limited to AIME’24 and AMC’23. Similar discrepancies are observed on MATH500, where different compute clusters yield different accuracy scores—even when model, seeds, and decoding parameters are held constant. This further emphasizes the need for hardware and software standardization when reporting benchmark results.
A.3 Prompt Variants and Template Settings
We provide the exact templates used for our three prompt settings in Table 5: Math, Default, and No Template. These formats are based on the DeepSeek tokenizer but adapted for each model’s specific chat template. Our results (in 3.4) indicate that instruction-tuned models are highly sensitive to prompt formatting, with performance degrading significantly when prompts deviate from their training-time structure.
@@ A.3 Prompt Variants and Template Settings
No Template | {Question} |
A.4 Effect of Output Length Limits
@@
A.5 Response Length vs. Accuracy — Per-Model Breakdown
@@
Figures 17 and 18 reveal that the overall trend observed in the main paper holds consistently across nearly all models: incorrect responses tend to be longer than correct ones.
These results reinforce the idea that excessively long outputs often indicate failure modes such as hallucinated reasoning, verbose overthinking, or degenerate loops. Importantly, this correlation persists well below the maximum sequence length, ruling out truncation as the sole cause.
Across all models, longer responses are a consistent marker of incorrect outputs, making response length a useful signal for detecting low-confidence or erroneous reasoning chains.

### v2

@@ 5.1 Are Incorrect Responses Longer?
Yes. We find the trend is consistent across both RL- and SFT-trained models (see Appendix figures 22 and 23 for detailed per-model breakdown ). We consistently observe that the effect is more pronounced in RL-trained models (displayed on the left) than in SFT-trained models (displayed on the right). As detailed in the Appendix, both the Qwen 2.5 Math base exhibit a slight shift in length, though this shift is notably more evident in R1-distill and subsequent RL-trained models.
@@
No, truncation is not the primary cause. While responses approaching the 32,000-token limit are almost always incorrect, the correlation between length and errors persists even for complete responses well below this limit (e.g., 10,000-15,000 tokens). This indicates that excessive length itself, rather than truncation, is associated with reasoning failures.
5.2 Is There Diversity Collapse in Reasoning Training?
Dang et al. (2025) and Yue et al. (2025a) have reported a counterintuitive phenomenon in reasoning models: improvements in Pass@1 achieved through supervised fine-tuning or RL can reduce Pass@k performance due to diminished output diversity—a phenomenon termed diversity collapse. Theoretical analyses attribute this collapse to the model concentrating too much probability mass on a single reasoning path, while current decoding strategies fail to recover the lost diversity.
To examine these claims, we compare the Pass@k performance (for ) of two RL-trained models (DeepScaleR-1.5B and FastCuRL-1.5B) against their base model DeepSeek-R1-Distill-Qwen-1.5B across all datasets. Figures 12 and 24 show the delta in Pass@k relative to DeepSeek-R1-Distill-Qwen-1.5B.
Findings. We do observe a minor diversity collapse. Gains in Pass@1 generally come with regression in Pass@k, though the magnitude of the decay varies.
5.3 Spurious Prompts and Qwen2.5-Math-7B
[…]
@@ 6 Conclusion
Our study shows that much of the perceived progress in LLM-based reasoning, particularly in mathematical benchmarks, rests on unstable and often non-reproducible foundations. We find that minor differences in sampling parameters, prompt formatting, hardware, and software configurations can lead to major shifts in reported performance—casting doubt on many recent empirical claims. Reinforcement learning methods, while promising in theory, offer at best modest gains in practice and are prone to overfitting, especially on small benchmarks like AIME’24. In contrast, supervised finetuning continues to deliver consistent, generalizable improvements across a wide range of benchmarks and model sizes. To address these challenges, we advocate for standardized, transparent evaluation protocols. Our open-sourced framework, complete with Dockerized environments, seed-averaged metrics, and robust answer matching, provides reproducible foundations for future research. We hope this work shifts the focus from leaderboard chasing to methodological rigor—ensuring that future claims of progress in reasoning are both meaningful and measurable.
@@ Acknowledgments
Appendix
We now provide thorough details about the benchmark, baselines, and prompts.
Appendix A Bootstrapping Results on Additional Datasets
To complement our analysis in Section 3, we present bootstrapped variance results on two additional datasets: AMC’23 and MATH500. As shown in Figures 14 and 15, high variance in Pass@1 persists even when averaging over multiple seeds (), mirroring the trends observed on AIME’24. These results reinforce our conclusion that small benchmark sizes yield unstable estimates and that robust performance reporting requires multiple seed runs. Even for larger Benchmarks, like MATH500 (Figure 16), Minerva (Figure 17) and Olympiad Bench (Figure 18) the estimates remain volative, yet there magnitude is much smaller. An overview of the exact variance values can be found in Table 5.
| Model | |||||||||
| AIME’24 | |||||||||
| DeepSeek-R1-Distill-1.5B | 40.044 | 19.841 | 13.269 | 7.762 | 3.754 | 2.409 | 1.686 | 1.054 | 0.365 |
| DeepScaleR-1.5B | 27.452 | 13.477 | 9.117 | 5.342 | 2.495 | 1.587 | 1.149 | 0.705 | 0.246 |
| FastCuRL-1.5B | 28.807 | 14.647 | 9.488 | 5.658 | 2.711 | 1.709 | 1.247 | 0.757 | 0.256 |
| AIME’25 | |||||||||
| DeepSeek-R1-Distill-1.5B | 24.300 | 11.954 | 7.759 | 4.568 | 2.199 | 1.419 | 1.017 | 0.628 | 0.209 |
| DeepScaleR-1.5B | 17.311 | 8.561 | 5.597 | 3.387 | 1.605 | 0.988 | 0.726 | 0.445 | 0.156 |
| FastCuRL-1.5B | 16.201 | 8.232 | 5.412 | 3.221 | 1.504 | 1.000 | 0.710 | 0.424 | 0.145 |
| AMC’23 | |||||||||
| DeepSeek-R1-Distill-1.5B | 22.544 | 11.241 | 7.301 | 4.288 | 2.093 | 1.315 | 0.942 | 0.585 | 0.202 |
| DeepScaleR-1.5B | 22.377 | 10.851 | 7.167 | 4.228 | 2.079 | 1.295 | 0.936 | 0.568 | 0.192 |
| FastCuRL-1.5B | 23.671 | 11.841 | 7.649 | 4.605 | 2.149 | 1.404 | 0.980 | 0.594 | 0.211 |
| MATH500 | |||||||||
| DeepSeek-R1-Distill-1.5B | 1.175 | 0.561 | 0.367 | 0.226 | 0.110 | 0.069 | 0.049 | 0.030 | 0.010 |
| DeepScaleR-1.5B | 1.050 | 0.513 | 0.351 | 0.198 | 0.099 | 0.062 | 0.044 | 0.027 | 0.009 |
| FastCuRL-1.5B | 0.939 | 0.447 | 0.305 | 0.179 | 0.085 | 0.055 | 0.040 | 0.024 | 0.008 |
| Minerva | |||||||||
| DeepSeek-R1-Distill-1.5B | 3.485 | 1.741 | 1.143 | 0.665 | 0.326 | 0.206 | 0.149 | 0.091 | 0.031 |
| DeepScaleR-1.5B | 2.385 | 1.179 | 0.769 | 0.452 | 0.220 | 0.143 | 0.102 | 0.061 | 0.022 |
| FastCuRL-1.5B | 2.703 | 1.347 | 0.876 | 0.527 | 0.244 | 0.161 | 0.113 | 0.069 | 0.024 |
| Olympiad Bench | |||||||||
| DeepSeek-R1-Distill-1.5B | 1.269 | 0.622 | 0.427 | 0.241 | 0.118 | 0.076 | 0.053 | 0.032 | 0.011 |
| DeepScaleR-1.5B | 1.018 | 0.505 | 0.334 | 0.195 | 0.093 | 0.059 | 0.043 | 0.026 | 0.009 |
| FastCuRL-1.5B | 0.908 | 0.463 | 0.305 | 0.179 | 0.086 | 0.054 | 0.039 | 0.024 | 0.008 |
Appendix B Prompt Variants and Template Settings
We provide the exact templates used for our three prompt settings in Table 6: Math, Default, and No Template. These formats are based on the DeepSeek tokenizer but adapted for each model’s specific chat template. Our results (in Section 3.3) indicate that instruction-tuned models are highly sensitive to prompt formatting, with performance degrading significantly when prompts deviate from their training-time structure.
@@ Appendix B Prompt Variants and Template Settings
[…]
| Prompt | Example |
[…]
Appendix C Hardware Differences
In Figure 19, we show that similar discrepancies are observed on MATH500 as observed in AIME and AMC – showing the variance from hardware is observed even on larger test sets.
Appendix D Effect of Output Length Limits
@@
Appendix E Response Length vs. Accuracy — Per-Model Breakdown
@@
Figures 22 and 23 reveal that the overall trend observed in the main paper holds consistently across nearly all models: incorrect responses tend to be longer than correct ones.
These results reinforce the idea that excessively long outputs often indicate failure modes such as hallucinated reasoning, verbose overthinking, or degenerate loops. Importantly, this correlation persists well below the maximum sequence length, ruling out truncation as the sole cause. Across all models, longer responses are a consistent marker of incorrect outputs, making response length a useful signal for detecting low-confidence or erroneous reasoning chains.
Appendix F Diversity Collapse
In addition to the diversity collapse discussed in Section 5.2 and Figure 12 the results for additional datasets are shown in Figure 24. For DeepSeek-R1-Distill-1.5B (SFT-trained) we could not replicate a diversity collapse as shown in Figure 25.
Appendix G Optimal Decoding Parameters
Empirically, the decoding parameters in Table 8 consistently produced optimal performance for the models we evaluated.
| Model Name | Temperature | Top-p |
| agentica-org/DeepScaleR-1.5B-Preview | 1.0 | 0.7 |
| bespokelabs/Bespoke-Stratos-7B | 1.0 | 0.9 |
| deepseek-ai/DeepSeek-R1-Distill-Qwen-1.5B | 0.9 | 0.7 |
| deepseek-ai/DeepSeek-R1-Distill-Qwen-7B | 0.9 | 0.7 |
| hkust-nlp/Qwen-2.5-7B-SimpleRL-Zoo | 0.5 | 0.5 |
| Intelligent-Internet/II-Thought-1.5B-Preview | 1.0 | 0.5 |
| knoveleng/Open-RS1 | 1.0 | 0.6 |
| knoveleng/Open-RS2 | 0.9 | 0.8 |
| knoveleng/Open-RS3 | 0.5 | 0.9 |
| l3lab/L1-Qwen-1.5B-Exact | 0.5 | 0.7 |
| l3lab/L1-Qwen-1.5B-Max | 0.7 | 0.9 |
| Nickyang/FastCuRL-1.5B-Preview | 1.0 | 0.7 |
| Open-Reasoner-Zero/Open-Reasoner-Zero-7B | 0.7 | 0.5 |
| open-thoughts/OpenThinker-7B | 0.8 | 0.95 |
| qihoo360/Light-R1-7B-DS | 0.7 | 1.0 |
| RUC-AIBOX/STILL-3-1.5B-preview | 1.0 | 0.6 |
| simplescaling/s1.1-7B | 1.0 | 0.9 |
| GAIR/LIMR | 0.6 | 0.6 |
| open-r1/OpenR1-Qwen-7B | 0.8 | 0.9 |
| Qwen/Qwen2.5-1.5B-Instruct | 0.2 | 1.0 |
| Qwen/Qwen2.5-7B-Instruct | 0.4 | 0.95 |
| Qwen/Qwen2.5-Math-1.5B | 0.7 | 0.5 |
| Qwen/Qwen2.5-Math-7B | 0.5 | 0.5 |
| sail/Qwen2.5-Math-1.5B-Oat-Zero | 0.6 | 0.6 |
| sail/Qwen2.5-Math-7B-Oat-Zero | 0.6 | 0.6 |